OpenStudy (anonymous):

Find the equation of a line given the point and slope below. (6, 6) no slope

OpenStudy (anonymous):

I figured it was y=x, but apparently not...

OpenStudy (anonymous):

I mean y = 6

OpenStudy (anonymous):

y=6 it is!...Not. Sorry.

OpenStudy (anonymous):

Maybe it's x=6, lemme check.

OpenStudy (anonymous):

Yeah it is. Thanks for your help though!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (2, 1) m = 0

OpenStudy (anonymous):

1 = 0(2) + b 1 = 0 + b 1 + b y = 0x +1

OpenStudy (anonymous):

It's a three-character answer...

OpenStudy (anonymous):

y = 1

OpenStudy (anonymous):

because 0x is nothing

OpenStudy (anonymous):

There it is! y = 1. Thanks again!

OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. for point (2, 1) , m = 0 y=mx+b substitute the point and the slope into the equation, then solve for b then write it back in the form of y=mx+b without the point

OpenStudy (anonymous):

Completeidiot, I already know how to do the equation. It was just a couple that I was confused on.

OpenStudy (anonymous):

slope of 0 mean a horizontal line no slope or infinite line means a vertical line

OpenStudy (anonymous):

Again, I know...